# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 3.3. Mumford construction, polytope version
We discuss now a polytopal version of the Mumford degeneration, which outputs a relatively projective degeneration, together with a relatively ample line bundle. Furthermore, it is isomorphic to the fan construction as in Section 3.1, for an appropriate choice of fan $\mathcal{S}$} in ${\bf N}_{\mathbb{R}}\times{\mathbb{R}}^{k}$. Our approach is, in part, inspired by Gross–Siebert [31, Sec. 2], and their construction of canonical theta functions, building on the classical theory of theta functions, see e.g. [46, Prop. II.1.3 and Thm. II.1.3]. It is primarily based on a “PL version” of the classical theory, in line with Alexeev–Nakamura [7].
@@
Consider the standard torus ${\mathbb{T}}^{g}\coloneqq{\mathbb{R}}^{g}/{\mathbb{Z}}^{g}={\bf M}_{\mathbb{R}}/{\bf M}$. We define ${\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L}$ to be the sheaf on ${\mathbb{T}}^{g}$ of $\mathbb{Z}$}-piecewise linear functions modulo the subsheaf of $\mathbb{Z}$}-linear functions. On an open set $U\subset{\bf M}_{\mathbb{R}}/{\bf M}$, the sections ${\mathbb{Z}}{\rm PL}(U)\coloneqq\{f\colon U\to{\mathbb{R}}\}$ consist of continuous, piecewise linear functions, which in a domain $D\subset U$ of linearity are of the form $f|_{D}({\bf m})=a_{1}m_{1}+\cdots+a_{g}m_{g}+a_{g+1}$ with $a_{j}\in{\mathbb{Z}}$. Here $m_{i}$ are integral coordinates on $\bf M$$_{\mathbb{R}}$ which define local coordinates on $D$. Similarly, ${\mathbb{Z}}{\rm L}$ is the sheaf of locally $\mathbb{Z}$}-linear functions on $U$, of the same shape. A section ${\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L}(U)$ can be understood globally on $U$ in terms of its “bending locus”, see Definition 3.16.
@@
Let $\{\overline{b}_{1},\dots,\overline{b}_{k}\}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ be a collection of convex sections, with nontrivial bending in every direction, i.e. for every ${\bf m}\in{\bf M}$, there exists some $i\in\{1,\dots,k\}$ for which $b_{i}(x+{\bf m})-b_{i}(x)\not\equiv 0$ is not identically zero. Equivalently, $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$ cuts ${\mathbb{T}}^{g}$ into polytopes. We say that the $\overline{b}_{1},\dots,\overline{b}_{k}$ are dicing if the polyhedral decomposition $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$ has integral vertices.
The dicing condition is quite restrictive, since only the origin of ${\mathbb{T}}^{g}$ may appear as a vertex of $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$. We will relax this hypothesis in Construction 3.37.
@@
where the coefficients $c_{\overline{v}_{1}\overline{v}_{2}}^{\overline{v}_{3}}(u_{1},\dots,u_{k})\in{\mathbb{Z}}[[u_{1},\dots,u_{k}]]$ are integral power series, as opposed to simply Laurent series, by the convexity of the $b_{i}$, see e.g. [31, Eqn. (2.5)]. Note that to get a nonzero coefficient, there must be a lift of $\overline{v}_{3}$ of the form
@@
That the generic fiber of (17) is an abelian variety, also follows from the classical theory of theta functions.
@@
To “twist” the construction, as in Constructions 3.3, 3.8, 3.10, by some continuous parameters $a=(a_{ij})\in{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$, and produce a universal degeneration which represents all possible continuous moduli of degenerations of the specified combinatorial type, we must introduce appropriate coefficients
@@
We say that $\{b_{1},\dots,b_{k}\}$ are $\tfrac{1}{d}$-dicing the corresponding overgraph $\Gamma$ of $\Gamma(b_{1},\dots,b_{k})\subset{\bf M}_{\mathbb{R}}\times{\mathbb{R}}^{k}$ is a $\tfrac{1}{d}({\bf M}\times{\mathbb{Z}}^{k})$-integral polyhedron.
Construction 3.25.
Like Construction 3.10 vis-à-vis Construction 3.8, we generalize Construction 3.20 to the case where $\{\overline{b}_{1},\dots,\overline{b}_{k}\}$ are the extremal rays of a convex polyhedral cone $\mathbbm{b}\subset H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ mapping isomorphically to a convex polyhedral cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$, under the map $\overline{b}_{i}\mapsto B_{i}$ with $B_{i}\in\operatorname{Sym}^{2}{\bf M}^{\vee}$ defined in (21). We replace $(b_{1},\dots,b_{k})$ by the PL function
@@
Otherwise, the details of Construction 3.20 are the same. The output is a relatively projective degeneration of abelian varieties
@@
over an analytic tubular neighborhood of the torus fixed point of $Y(\mathbbm{b})$. Performing this construction with the universal twist by $a\in{\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle\otimes{\mathbb{C}}^{*}$, for ${\mathbb{B}}={\mathbb{R}}_{\geq 0}\{B_{1},\dots,B_{k}\}$, we may produce a relatively projective analytic extension of the universal family
$X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{+}\subset\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}},$
see Proposition 3.12. We will see in Proposition 6.12 that $X^{\rm univ}(\mathbbm{b})$ is not just relatively projective in the analytic category, but is the restriction of a projective morphism of quasiprojective varieties, in the algebraic category, to the analytic open subset $\widetilde{{\mathcal{A}}}_{g}^{+}$. $\clubsuit$
@@
Our primary case of interest in [27] is where $\mathbbm{b}$ defines a simplicial cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$ and in this setting, Constructions 3.8 and 3.20 over a polydisk will suffice.
The next proposition follows directly from Construction 3.20:
@@ 3.4. Comparison of polytope and fan constructions
$\Theta_{\overline{v}}(z,u)=\sum_{{\bf m}\in{\bf M}}(z,u)^{w({\bf m}\cdot(v,\vec{r}))}$
@@ 3.5. Examples
Then $\Gamma=\{(m,\tfrac{1}{2}(m^{2}-m))\,:\,m\in{\mathbb{Z}}={\bf M}\}\subset{\bf M}\times{\mathbb{Z}}={\mathbb{Z}}^{2}$ and $\Gamma$ is the shaded region in Figure 2. The theta functions of weight $1$, 22, and $3$ are
@@
This example is originally attributed to Deligne [44, Sec. 7]; called by Mumford the “keystone” of the compactification of ${\mathcal{A}}_{2}$. The figure is similar to Figure 6, but we now use three different colors, to indicate the different bending loci ${\rm Bend}(b_{i})$, for $i=1,2,3$. See Figure 9.
@@
To understand the fibers of $X(b_{1},b_{2},b_{3})=X\to\Delta^{3}$ over the various coordinate subspaces, we refer to Figure 10. By Proposition 3.27, the polytopes of the components of the fiber over $(u_{1},u_{2},u_{3})\in\Delta^{3}$ of the Mumford construction can be read off from the the bending loci of the $b_{i}$ for which $V(u_{i})=0$. The fiber over the origin is the union
@@ 3.6. Base change and Veronese embedding
Let $X(b_{1},\dots,b_{k})\to\Delta^{k}$ be a Mumford construction on a collection of convex sections $b_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$, see Construction 3.20, and consider a monomial base change, of the form $\Delta^{n}\to\Delta^{k}$, such that the pullback of coordinates $u_{i}$ are of the form
@@
The base change of the Mumford degeneration $X(b_{1},\dots,b_{k})\to\Delta^{k}$ along a monomial base change $\Delta^{n}\to\Delta^{k}$ is the Mumford construction $X(c_{1},\dots,c_{n})\to\Delta^{n}$ associated to the convex PL functions
@@
A simple case is exhibited by Examples 3.33, 3.34, where we make the monomial base change $\Delta\to\Delta^{3}$, $w\mapsto(w^{r_{1}},w^{r_{2}},w^{r_{3}})$ to Example 3.34, to get Example 3.33.
@@
consisting of the theta functions of weights $w$ divisible by $d$. Thus, the Veronese subring may be constructed via a mild generalization of Construction 3.20, where we consider instead sections of ${\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L}$ on the real torus
${\mathbb{T}}^{g}(d)\coloneqq{\bf M}_{\mathbb{R}}/(d{\bf M})^{g}$
of side length $d$. It is simpler to rescale ${\mathbb{T}}^{g}(d)$ back to ${\mathbb{T}}^{g}$ and instead consider sections of $\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L}$ (this perspective also has the advantage that the functions $b_{i}$ are unchanged by taking the Veronese embedding). The passage to the Veronese subring suggests also a natural generalization, which allows us to relax the restrictive dicing condition (Def. 3.17):
Construction 3.37.
We further generalize Construction 3.20 to allow for convex sections
@@
for some positive integer $d$. Then, we may define theta functions similarly to formula (15), but only for the weights $w$ divisible by $d$. Assuming that the $\overline{b}_{i}$ are $\tfrac{1}{d}$-dicing on ${\mathbb{T}}^{g}$ (Def. 3.24), we may deduce from Remark 3.23 that $dL$ lifts to an ample line bundle on the resulting degeneration, which we denote by
@@
While the general fiber of the degeneration still admits a principal polarization $L$, only $dL$ extends to a line bundle on the total space, in general. Furthermore, the general fiber of the Mumford construction is the quotient by $d{\bf M}$, rather than $\bf M$}, of $\bf N$$\otimes{\mathbb{C}}^{*}$. $\clubsuit$
@@ 4.1. Matroids, graphs, and quadratic forms
By a theorem of Tutte [61], every regular matroid can be defined by a totally unimodular matrix, that is, a matrix all of whose minors (in particular, all entries) have determinant in $\{\pm 1,0\}$. Then, an integral realization of the matroid arises by considering the set of column vectors. More generally, any unimodular matrix—an integer entry matrix whose maximal minors have determinant in $\{\pm 1,0\}$—defines an integral realization of a regular matroid [63, Ch. 3, Thm. 3.1.1]. Equivalently, the lattice spanned by any collection of columns is saturated.

### v3

@@ 3.3. Mumford construction, polytope version
We discuss now a polytopal version of the Mumford degeneration, which outputs a relatively projective degeneration, together with a relatively ample line bundle. Furthermore, it is isomorphic to the fan construction as in Section 3.1, for an appropriate choice of fan $\mathcal{S}$} in ${\bf N}_{\mathbb{R}}\times{\mathbb{R}}^{k}$. Our approach is, in part, inspired by Gross–Siebert [32, Sec. 2], and their construction of canonical theta functions, building on the classical theory of theta functions, see e.g. [48, Prop. II.1.3 and Thm. II.1.3]. It is primarily based on a “PL version” of the classical theory, in line with Alexeev–Nakamura [7].
@@
Consider the standard torus ${\mathbb{T}}^{g}\coloneqq{\mathbb{R}}^{g}/{\mathbb{Z}}^{g}={\bf M}_{\mathbb{R}}/{\bf M}$. We define ${\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L}$ to be the sheaf on ${\mathbb{T}}^{g}$ of $\mathbb{Z}$}-piecewise linear functions modulo the subsheaf of $\mathbb{Z}$}-linear functions. On an open set $U\subset{\bf M}_{\mathbb{R}}/{\bf M}$, the sections ${\mathbb{Z}}{\rm PL}(U)\coloneqq\{f\colon U\to{\mathbb{R}}\}$ consist of continuous, piecewise linear functions, which in a domain $D\subset U$ of linearity are of the form
(18) $f|_{D}({\bf m})=a_{1}m_{1}+\cdots+a_{g}m_{g}+a_{g+1}$
with $a_{j}\in{\mathbb{Z}}$. Here $m_{i}$ are integral coordinates on $\bf M$$_{\mathbb{R}}$ which define local coordinates on $D$. Similarly, ${\mathbb{Z}}{\rm L}$ is the sheaf of locally $\mathbb{Z}$}-linear functions on $U$, of the same shape. A section ${\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L}(U)$ can be understood globally on $U$ in terms of its “bending locus” (Definition 3.17).
@@
Let $\{\overline{b}_{1},\dots,\overline{b}_{k}\}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ be a collection of convex sections, with nontrivial bending in every direction, i.e. for every non-zero ${\bf m}\in{\bf M}$, there exists some $i\in\{1,\dots,k\}$ for which $b_{i}(x+{\bf m})-b_{i}(x)\not\equiv 0$ is not identically zero. Equivalently, $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$ cuts ${\mathbb{T}}^{g}$ into polytopes. We say that the $\overline{b}_{1},\dots,\overline{b}_{k}$ are dicing if the polyhedral decomposition $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$ has integral vertices.
The dicing condition is quite restrictive, since only the origin of ${\mathbb{T}}^{g}$ may appear as a vertex of $\bigcup_{i}{\rm Bend}(\overline{b}_{i})$. We will relax this hypothesis in Construction 3.38.
@@
where the coefficients $c_{\overline{v}_{1}\overline{v}_{2}}^{\overline{v}_{3}}(u_{1},\dots,u_{k})\in{\mathbb{Z}}[[u_{1},\dots,u_{k}]]$ are integral power series, as opposed to simply Laurent series, by the convexity of the $b_{i}$, see e.g. [32, Eqn. (2.5)]. Note that to get a nonzero coefficient, there must be a lift of $\overline{v}_{3}$ of the form
@@
That the generic fiber of (21) is an abelian variety also follows from the classical theory of theta functions.
@@
To “twist” the construction, as in Constructions 3.3, 3.8, 3.11, by some continuous parameters $a=(a_{ij})\in{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$, and produce a universal degeneration which represents all possible continuous moduli of degenerations of the specified combinatorial type, we must introduce appropriate coefficients
@@
We say that $\{b_{1},\dots,b_{k}\}$ are $\tfrac{1}{d}$-dicing if they are $\mathbb{Q}$}-piecewise linear, the corresponding overgraph $\Gamma$ of $\Gamma(b_{1},\dots,b_{k})\subset{\bf M}_{\mathbb{R}}\times{\mathbb{R}}^{k}$ is a $\tfrac{1}{d}({\bf M}\times{\mathbb{Z}}^{k})$-integral polyhedron, and in the local form (18), the slopes $a_{1},\dots,a_{g}\in{\mathbb{Z}}$ are still integral, but we allow $a_{g+1}\in\tfrac{1}{d}{\mathbb{Z}}$. We denote sheaves of functions with such a local form by $\tfrac{1}{d}{\mathbb{Z}}{\rm PL}$ and $\tfrac{1}{d}{\mathbb{Z}}{\rm L}$.
Construction 3.26.
Like Construction 3.11 vis-à-vis Construction 3.8, we generalize Construction 3.21 to the case where $\{\overline{b}_{1},\dots,\overline{b}_{k}\}$ are the extremal rays of a convex polyhedral cone $\mathbbm{b}\subset H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$ mapping isomorphically to a convex polyhedral cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$ under the map $\overline{b}_{i}\mapsto B_{i}$ with $B_{i}\in\operatorname{Sym}^{2}{\bf M}^{\vee}$ defined in (25). We replace $(b_{1},\dots,b_{k})$ by the PL function
@@
Otherwise, the details of Construction 3.21 are the same. The output is a relatively projective degeneration of abelian varieties
@@
over an analytic tubular neighborhood of the torus fixed point of $Y(\mathbbm{b})$. Performing this construction with the universal twist by $a\in{\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle\otimes{\mathbb{C}}^{*}$, for ${\mathbb{B}}={\mathbb{R}}_{\geq 0}\{B_{1},\dots,B_{k}\}$, and extending/descending over the toroidal extension $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ as in Construction 3.11, we may produce an analytically-locally relatively projective, analytic extension of the universal family
$X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}},$
see Proposition 3.13 (we descend from an open set $T_{\rm max}(\mathbbm{b})\subset Y(\mathbbm{b})$ as in Construction 3.11 to verify the analytic-local projectivity). We will show in Section 6 that $X^{\rm univ}(\mathbbm{b})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ is an étale-locally projective morphism of algebraic spaces over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. $\clubsuit$
@@
Our primary case of interest in [27] is where $\mathbbm{b}$ defines a simplicial cone ${\mathbb{B}}\subset{\mathcal{P}}_{g}^{+}$ and in this setting, Constructions 3.8 and 3.21 over a polydisk will suffice.
The next proposition follows directly from Construction 3.21:
@@ 3.4. Comparison of polytope and fan constructions
(26) $\Theta_{\overline{v}}(z,u)=\sum_{{\bf m}\in{\bf M}}(z,u)^{w({\bf m}\cdot(v,\vec{r}))}$
@@ 3.5. Examples
Then $\Gamma$ is the shaded region in Figure 2, and the convex hull of $\{(m,\tfrac{1}{2}(m^{2}-m)):m\in{\bf M}\}\subset{\bf M}\times{\mathbb{Z}}={\mathbb{Z}}^{2}$. The theta functions of weight $1$, 22, and $3$ are
@@
as only $0/3\in\tfrac{1}{3}{\bf M}/{\bf M}\subset{\mathbb{R}}/{\mathbb{Z}}$ lies in multiple domains of linearity of $\overline{b}$.
@@
This example is originally attributed to Deligne [46, Sec. 7]; called by Mumford the “keystone” of the compactification of ${\mathcal{A}}_{2}$. The figure is similar to Figure 6, but we now use three different colors, to indicate the different bending loci ${\rm Bend}(b_{i})$, for $i=1,2,3$. See Figure 9.
@@
To understand the fibers of $X(b_{1},b_{2},b_{3})=X\to\Delta^{3}$ over the various coordinate subspaces, we refer to Figure 10. By Proposition 3.28, the polytopes of the components of the fiber over $(u_{1},u_{2},u_{3})\in\Delta^{3}$ of the Mumford construction can be read off from the bending loci of the $b_{i}$ for which $V(u_{i})=0$. The fiber over the origin is the union
@@ 3.6. Base change and Veronese embedding
Let $X(b_{1},\dots,b_{k})\to\Delta^{k}$ be a Mumford construction on a collection of convex sections $b_{i}\in H^{0}({\mathbb{T}}^{g},{\mathbb{Z}}{\rm PL}/{\mathbb{Z}}{\rm L})$, see Construction 3.21, and consider a monomial base change, of the form $\Delta^{n}\to\Delta^{k}$, such that the pullback of coordinates $u_{i}$ are of the form
@@
The base change of the Mumford degeneration $X(b_{1},\dots,b_{k})\to\Delta^{k}$ along a monomial morphism $\Delta^{n}\to\Delta^{k}$ is the Mumford construction $X(c_{1},\dots,c_{n})\to\Delta^{n}$ associated to the convex PL functions
@@
A simple case is exhibited by Examples 3.34, 3.35, where we make the monomial base change $\Delta\to\Delta^{3}$, $w\mapsto(w^{r_{1}},w^{r_{2}},w^{r_{3}})$ to Example 3.35, to get Example 3.34.
@@
consisting of the theta functions of weights $w$ divisible by $d$. The passage to the Veronese subring suggests also a natural generalization of Construction 3.21, which allows us to relax the restrictive dicing condition:
Construction 3.38.
Consider convex sections
@@
for some positive integer $d$, see Definition 3.18. Then, we may define theta functions similarly to formula (19), but only for the weights $w$ divisible by $d$. Assuming that the $\overline{b}_{i}$ are $\tfrac{1}{d}$-dicing on ${\mathbb{T}}^{g}$ (Def. 3.25), we may deduce from Remark 3.24 that $dL$ lifts to an ample line bundle on the resulting degeneration, which we denote by
@@
While the general fiber of the degeneration still admits a principal polarization $L$, only $dL$ extends to a line bundle on the total space, in general.
In terms of the polytope, we still take the overgraph $\Gamma=\Gamma(b_{1},\dots,b_{k})+({\mathbb{R}}_{\geq 0})^{k}$, which is now only a $\tfrac{1}{d}({\bf M}\times{\mathbb{Z}}^{k})$-integral polyhedron. Then, as in Remark 2.21, we consider ${\rm Cone}(\Gamma)\subset({\bf M}_{\mathbb{R}}\times{\mathbb{R}}^{k})\times{\mathbb{R}}$ but one only considers monomials, and their $\bf M$}-averagings (26) to theta functions, lying in $({\bf M}\times{\mathbb{Z}}^{k})\times d{\mathbb{Z}}$. The general fiber of the Mumford construction is still the quotient by $\bf M$} of $\bf N$$\otimes{\mathbb{C}}^{*}$; in particular, the exact sequence (11) still holds. $\clubsuit$
Remark 3.39.
The isomorphism type of the degeneration $X(d\mid b_{1},\dots,b_{k})\to\Delta^{k}$ does not depend on the lifts $b_{i}$ of $\overline{b}_{i}$ but the choice of origin section does, since different lifts of $\overline{b}_{i}$ shift the normal fan, and thus affect which subtorus forms the origin section, see Remark 3.9. The same applies to Construction 3.26—to produce an extension of the universal family over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ requires a choice of lift of cone $\mathbbm{b}\subset H^{0}({\mathbb{T}}^{g},\tfrac{1}{d}{\mathbb{Z}}{\rm PL}/\tfrac{1}{d}{\mathbb{Z}}{\rm L})$ into $H^{0}({\bf M}_{\mathbb{R}},\tfrac{1}{d}{\mathbb{Z}}{\rm PL})$ as the gluing with the universal family $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$ depends on a choice of origin section.
@@ 4.1. Matroids, graphs, and quadratic forms
By a theorem of Tutte [65], every regular matroid can be defined by a totally unimodular matrix, that is, a matrix all of whose minors (in particular, all entries) have determinant in $\{\pm 1,0\}$. Then, an integral realization of the matroid arises by considering the set of column vectors. More generally, any unimodular matrix—an integer entry matrix whose maximal minors have determinant in $\{\pm 1,0\}$—defines an integral realization of a regular matroid [68, Ch. 3, Thm. 3.1.1]. Equivalently, the lattice spanned by any collection of columns is saturated.